$SPAD/src/input exlimit.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool exlimit.output
)set message test on
)set message auto off
)clear all
--S 1 of 13
limit((x**2 - 3*x + 2)/(x**2 - 1),x = 1)
--R
--R
--R          1
--R   (1)  - -
--R          2
--R            Type: Union(OrderedCompletion(Fraction(Polynomial(Integer))),...)
--E 1
)clear all
--S 2 of 13
complexLimit((2 + z)/(1 - z),z = %infinity)
--R
--R
--R   (1)  - 1
--R                      Type: OnePointCompletion(Fraction(Polynomial(Integer)))
--E 2
--S 3 of 13
limit(sin(x)/x,x = %plusInfinity)
--R
--R
--R   (2)  0
--R                      Type: Union(OrderedCompletion(Expression(Integer)),...)
--E 3
--S 4 of 13
complexLimit(sin(x)/x,x = %infinity)
--R
--R
--R   (3)  "failed"
--R                                                    Type: Union("failed",...)

--E 4
)clear all
--S 5 of 13
limit(x * log(x),x = 0,"right")
--R
--R
--R   (1)  0
--R                      Type: Union(OrderedCompletion(Expression(Integer)),...)
--E 5
--S 6 of 13
limit(x * log(x),x = 0)
--R
--R
--R   (2)  [leftHandLimit= "failed",rightHandLimit= 0]
--IType: Union(Record(leftHandLimit: Union(OrderedCompletion(...
--E 6
)clear all
--S 7 of 13
limit(sqrt(y**2)/y,y = 0)
--R
--R
--R   (1)  [leftHandLimit= - 1,rightHandLimit= 1]
--IType: Union(Record(leftHandLimit: Union(OrderedCompletion(...
--E 7
--S 8 of 13
limit(sqrt(1 - cos(t))/t,t = 0)
--R
--R
--R                            1                    1
--R   (2)  [leftHandLimit= - ----,rightHandLimit= ----]
--R                           +-+                  +-+
--R                          \|2                  \|2
--IType: Union(Record(leftHandLimit: Union(OrderedCompletion(...
--E 8
)clear all
--S 9 of 13
limit(sqrt(3*x**2 + 1)/(5*x),x = %plusInfinity)
--R
--R
--R         +-+
--R        \|3
--R   (1)  ----

--R          5
--R                      Type: Union(OrderedCompletion(Expression(Integer)),...)
--E 9
--S 10 of 13
limit(sqrt(3*x**2 + 1)/(5*x),x = %minusInfinity)
--R
--R
--R           +-+
--R          \|3
--R   (2)  - ----
--R            5
--R                      Type: Union(OrderedCompletion(Expression(Integer)),...)
--E 10
)clear all
--S 11 of 13
limit(sinh(a*x)/tan(b*x),x = 0)
--R
--R
--R        a
--R   (1)  -
--R        b
--R                      Type: Union(OrderedCompletion(Expression(Integer)),...)
--E 11
)clear all
--S 12 of 13
limit(z * sin(1/z),z = 0)
--R
--R
--R   (1)  0
--R                      Type: Union(OrderedCompletion(Expression(Integer)),...)
--E 12
--S 13 of 13
complexLimit(z * sin(1/z),z = 0)
--R
--R
--R   (2)  "failed"
--R                                                    Type: Union("failed",...)
--E 13
)spool
)lisp (bye)

———-

References^{[1] nothing}

6